OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form.

OpenStudy (anonymous):

Use point-slope form and simplify from there. \[y - y_{1} = m(x - x_{1})\]

OpenStudy (anonymous):

Standard form is Ax + By = C

OpenStudy (anonymous):

I know the equations...

OpenStudy (anonymous):

So just use them then.

OpenStudy (anonymous):

I got 4x + 4y = 6, is that right?

OpenStudy (anonymous):

I got x - 4y = -6

OpenStudy (anonymous):

\[m = \frac{3 - 2}{6 - 2} = \frac{1}{4}\]\[y - y_{1} = m(x - x_{1})\]\[y - 2 = \frac{1}{4}(x - 2)\]\[y - 2 = \frac{x}{4} -\frac{1}{2}\]\[4(y - 2 = \frac{x}{4} -\frac{1}{2})\]\[4y - 8 = x - 2\]\[x - 4y = -6\]